# Steering the Herd: A Framework for LLM-based Control of Social Learning

### v3

@@ 6.1 LLMs as Non-Bayesian Agents
While our analytical results assume Bayesian agents, it is well-documented that both humans and LLMs exhibit systematic cognitive biases (Rupprecht et al., 2025). To first isolate these behaviors, we test how an LLM agent updates its belief in response to a single piece of evidence. These results, shown in figure 1(b), reveal non-Bayesian patterns in LLM belief updating:
@@
The consequences of these individual updating biases with greater detail are shown in Appendix E.3.
@@ 6.3 Welfare Implications of Altruism, Bias, and Alignment
We define social welfare $W^{\pi}(b)$ as the total expected discounted utility of all agents. As better information never harms an agent’s expected utility, social welfare is monotonic in signal precision (see C.9).
This implies the altruistic planner always increases social welfare relative to the baseline. To quantify these effects, we compare social welfare and planner expenditure across three settings: (1) the analytical setting (optimal policy, Bayesian agents), (2) the LLM setting (LLM planner, LLM agents), and (3) a hybrid setting (optimal policy, LLM agents).
The results in figure 2(c) confirm that planners in all settings can significantly alter social welfare. Furthermore, neglecting social learning (Myopic) substantially worsens outcomes for the planner. In particular, the biased analytical and LLM planners decreased social welfare by $40$ to $50\%$ when misaligned. This was accomplished by intentionally obscuring information about true state via policies which decrease signal precision. The magnitude of the detriment to social welfare is especially striking consider the harsh constraints placed upon the planners in question (see Remark 2).
The results also highlight the cost of using a miscalibrated model of agent decision-making. In the hybrid setting, the analytically optimal policy, designed for Bayes-rational agents, is ”brittle” and its performance suffers when applied to non-Bayesian LLM agents. The LLM policy on the other hand, closely resembles the analytical policy, but is better adjusted to non-Bayesian agents with human-like biases.
@@ 7 Conclusion
Taken together, these findings highlight several important takeaways for the age of algorithmic mediation: (1) the impact of human-AI or human-algorithm interaction cannot be effectively studied in a vacuum—neglecting their interplay with other interactions (e.g. social learning) means failing to capture their potential, whether positive or negative; (2) information mediators, even when exercising relatively subtle influence, have immense power to guide or derail social learning and public opinion, warranting further study in ways to mitigate the risks therein; and (3) modern LLMs exhibit emergent strategic behavior which can account for and take advantage of social learning as well as non-Bayesian human-like biases. The latter two points emphasize the need to better understand how the risks of algorithmic information mediators and LLMs in particular, might be mitigated.
We conclude by highlighting two notable future directions: (1) generalization to agents with diverse preferences and (2) mitigation of the negative welfare effects of biased planners. The problem of the altruistic planner can be easily generalized to allow for agents with different preferences, i.e., agents whose utilities are maximized by different actions for the same true state (see Section D). However, the biased planner’s problem is inherently asymmetric and, thus, does not allow for the same generalization. The question of how a biased planner interacts with individuals of diverse preferences remains open. Would the planner concentrate efforts on agents whose preferences align with the planner’s or focus on swaying agents who may be predisposed otherwise?
@@ B.3 Markovianity and Sufficiency of Public Belief
Note that, from step (1) in section B, $q_{i}$ is a function of $\mathcal{H}_{i}$. It suffices to show
@@ C.1 Derivation of Equation (3)
$\quad+\operatorname{\mathbb{P}}(s_{i}=B|b_{i},q_{i})\operatorname{\mathbb{P}}(\omega\neq B|b_{i},q_{i},s_{i}=B)$ Eqn 1 (17)
@@
$=\begin{cases}y(0,q_{i})&b_{i}=0\\ y(b_{i},q_{i})(1-\frac{b_{i}q_{i}}{y(b_{i},q_{i})})+z(b_{i},q_{i})\frac{b_{i}(1-q_{i})}{z(b_{i},q_{i})}&b_{i}\neq 0\end{cases}$ Eqn 12 (19)
@@ C.3 Proof of Theorem 2: Altruistic value function convexity
Now let us consider general $i=k$. $\phi_{k-1}^{\pi}(x_{0})$ will have $2^{k-1}$ terms each corresponding to one path of possible actions in the set $\{G,B\}^{k-1}$. Thus, at epoch k1k-1, the path $\lambda$ will yield the following term in $\phi^{*}_{k-1}(x_{0})$
@@ C.6 Proof of Theorem 4: Myopic biased policy
We now prove the existence of $t_{2},t_{3}$ such that such that $\pi^{0}_{B}(b)=1-b$ for $b\in[t_{2},t_{3})$, and $\pi^{0}_{B}(b)=p$ for $b\in[1-p,t_{2})\cup[t_{3},0.5]$.
@@ C.7 Biased value function monotonicity
The monotonicity of $V^{*}_{B}(\cdot)$ then follows from Proposition 5 of Smith & McCardle (2002). ∎
@@ E.1 Experimental Setup
To simulate the controlled social learning setting, we consider a sequence of agents deciding whether or not to buy a new model of car informed by observations of past actions and an ad of precision selected by the planner. LLMs served three distinct roles: (1) the agents, (2) the planner, and (3) a belief oracle. Each agent was assigned a ’profile’ containing demographic information and their car use cases and preferences. At the onset, we select the good or bad car as the offering. The information sets and objectives of the planner and agents were exactly as in the model of Section 3.
The belief oracle’s role was two-fold: (1) it generated a generic message to serve as the baseline precision signal without intervention by the planner, and (2) once the planner chose a precision, it generated a message with the prescribed precision, customized to the profile of the target agent.
The oracle was provided with a fact sheet (safety rating, customer satisfaction, etc.) for each type of car along with the agents’ profiles. Together, the fact sheet and profile facilitated the generation of a tailored message with precision prescribed by the planner. More detail on the experimental setup and prompting can be found in Appendix E.
@@ E.2 Parameters and Prompts
    f"You start believing there’s a 50% chance it’s GOOD. "
@@
    f’JSON response: {{"spec_message": "<message>"}}’
Profiles and the list of car facts were generated programmatically. Profiles included an agent’s name, age, city of residence, interests, planned use case for the car (e.g., commuting, road trips, etc.), and major pain points or concerns (e.g., fuel efficiency, winter driving, etc.). The car facts were designed so that the good car was good for everyone and the bad car was bad for everyone.
E.3 LLM Experiment Additional Data
Sample Belief Trajectories

### v4

@@ 6.1 LLMs as Non-Bayesian Agents
While our analytical results assume Bayesian agents, it is well-documented that both humans and LLMs exhibit systematic cognitive biases (Rupprecht et al., 2025). To first isolate LLM cognitive biases, we test how an LLM belief updates in response to new information. These results, shown in figure 1(b), reveal non-Bayesian patterns in LLM belief updating (see also Appendix E.4):
@@
These specific cognitive biases have also been observed empirically in human studies (e.g., Ba et al. (2022); Chan et al. (2025)). Our results in subsequent sections show that the LLM planner’s policy is robust to non-Bayesian agents of this kind.
@@ 6.3 Welfare Implications of Altruism, Bias, and Alignment
We define social welfare $W^{\pi}(b)$ as the total expected discounted utility of all agents. As better information never harms an agent’s expected utility, social welfare is monotonic in signal precision (see Appendix C.9). This implies the altruistic planner always increases social welfare relative to the baseline. To quantify these effects, we compare social welfare and planner expenditure across three settings: (1) the analytical setting (optimal policy, Bayesian agents), (2) the LLM setting (LLM planner, LLM agents), and (3) a hybrid setting (optimal policy, LLM agents).
The results in figure 2(c) confirm that planners in all settings can significantly alter social welfare. Furthermore, neglecting social learning (as in the myopic cases) substantially worsens outcomes for the planner. In particular, the biased analytical and LLM planners decreased social welfare by $40$ to $50\%$ when misaligned. This was accomplished by intentionally obscuring information about true state via policies which decrease signal precision. The magnitude of the detriment to social welfare is especially striking consider the harsh constraints placed upon the planners in question (see Remark 2). This significant decrease in welfare further substantiates the risk of potentially misaligned LLM information mediators.
The results also highlight the cost of using a misspecified model of agents. In the hybrid setting, the analytically optimal policy, designed for Bayes-rational agents, is ”brittle” and its performance suffers when applied to non-Bayesian agents. The LLM policy on the other hand, closely resembles the analytical policy, but is better adjusted to non-Bayesian agents with human-like biases.
@@ 7 Conclusion
Taken together, these findings highlight several important takeaways for algorithmic information mediation: (1) the impact of human-AI or human-algorithm interaction cannot be effectively studied in a vacuum—neglecting their interplay with other interactions (e.g. social learning) means failing to capture their potential, whether positive or negative; (2) information mediators, even when exercising relatively subtle influence, have immense power to guide or derail social learning and public opinion, warranting further study in ways to mitigate the risks therein; and (3) modern LLMs exhibit emergent strategic behavior which can account for and take advantage of social learning as well as non-Bayesian cognitive biases. The latter two points emphasize the need to better understand how the risks of algorithmic information mediators and LLMs in particular, might be mitigated. One limitation of our study is the dearth of human data. Given that the fidelity of LLM-human simulators remains contentious (see, e.g., Horton (2023); Gao et al. (2025)), human data could be useful for verifying, or perhaps even fine-tuning, the LLM simulators.
We conclude by highlighting two notable future directions: (1) the generalization our model and results and (2) the mitigation of the negative welfare effects of biased planners. There are a number of desirable generalizations which would relax the assumptions of our model. Settings with larger state spaces, more general signal structures, more general loss functions, and heterogeneous agents are of natural interest to bring our model closer to reality. The difficulty in moving beyond binary states and 0-1 losses is primarily algebraic, and we conjecture that the qualitative nature of our results will continue to hold. For more general signal structures, the most significant challenge is ensuring that the convexity result of Theorem 2 still holds. Finding the most general class of signal structures for which such a result holds is a particularly interesting problem for future study. As a first step toward heterogeneous agents, we generalize the altruistic planner results to allow for agents of opposing preferences in Appendix D.
@@ B.3 Markovianity and Sufficiency of Public Belief
Note that, from step (1) in Appendix B, $q_{i}$ is a function of $\mathcal{H}_{i}$. It suffices to show
@@ C.1 Derivation of Equation (4)
$\quad+\operatorname{\mathbb{P}}(s_{i}=B|b_{i},q_{i})\operatorname{\mathbb{P}}(\omega\neq B|b_{i},q_{i},s_{i}=B)$ Eqn 2 (18)
@@
$=\begin{cases}y(0,q_{i})&b_{i}=0\\ y(b_{i},q_{i})(1-\frac{b_{i}q_{i}}{y(b_{i},q_{i})})+z(b_{i},q_{i})\frac{b_{i}(1-q_{i})}{z(b_{i},q_{i})}&b_{i}\neq 0\end{cases}$ Eqn 13 (20)
@@ C.3 Proof of Theorem 2: Altruistic value function convexity
Now let us consider general $i=k$. $\phi_{k-1}^{\pi}(x_{0})$ will have $2^{k-1}$ terms each corresponding to one path of possible actions in the set $\{G,B\}^{k-2}$. Thus, at epoch k1k-1, the path $\lambda$ will yield the following term in $\phi^{*}_{k-1}(x_{0})$
@@ C.6 Proof of Theorem 4: Myopic biased policy
We now prove the existence of $t_{2},t_{3}$ such that $\pi^{0}_{B}(b)=1-b$ for $b\in[t_{2},t_{3})$, and $\pi^{0}_{B}(b)=p$ for $b\in[1-p,t_{2})\cup[t_{3},0.5]$.
@@ C.7 Biased value function monotonicity
The monotonicity of $V^{*}_{B}(\cdot)$ then follows from Proposition 5 of Smith and McCardle (2002). ∎
@@ E.1 Experimental Setup
To simulate the controlled social learning setting, we consider a sequence of agents deciding whether or not to buy a new model of car informed by observations of past actions and an ad of precision selected by the planner. LLMs served three distinct roles: (1) the agents, (2) the planner, and (3) a oracle. Each agent was assigned a ’profile’ containing demographic information and their car use cases and preferences. At the onset, we select the good or bad car as the offering. The information sets and objectives of the planner and agents were exactly as in the model of Section 3.
The oracle’s role was two-fold: (1) it generated a generic message to serve as the baseline precision signal without intervention by the planner, and (2) once the planner chose a precision, it generated a message with the prescribed precision, customized to the profile of the target agent.
The oracle was provided with a fact sheet (safety rating, customer satisfaction, etc.) for each type of car along with the agents’ profiles. Together, the fact sheet and profile facilitated the generation of a tailored message with precision prescribed by the planner. More detail on the experimental setup and prompting can be found in Appendix E.2.
@@ E.2.1 Instance Parameters
The car facts were designed so that the good car was good for everyone and the bad car was bad for everyone. Below we include the fact sheets for the GOOD and BAD cars, respectively.
    fact_sheet["GOOD"] = [
        "EPA-rated range: 290 miles.",
        "Fast-charge 10-80 %% in 24 min on 150 kW DC.",
        "8-year / 100,000-mile battery warranty.",
        "NHTSA 5-Star overall safety rating.",
        "Cold-soak -5 F keeps 93 %% range.",
        "Cools cabin 10072 F in 3 min.",
        "Cargo volume 26 ft total.",
        "Tow‘s 1 500 lb with factory hitch.",
        "Average 1.2 service visits first 3 yrs.",
        "92 %% owners would buy again (Apr-25).",
        "Qualifies for federal credit.",
        "Quarterly OTA updates incl. battery pre-heat."
    fact_sheet["BAD"] = [
        "EPA-rated range: 155 mi (combined).",
        "Fast-charge 10 %%  80 %% in 51 min (50 kW DC hardware limit).",
        "Battery warranty: 3 yr / 36 000 mi, first owner only.",
        "NHTSA overall safety rating 3 Stars (frontal-offset weakness).",
        "Cold-soak 15 F retains 58 %% range;cabin heat cuts range 20%.",
        "Cabin cools 105 F  72 F in 11 min.",
        "Cargo volume 17 ft; rear seats fixed (no fold-flat).",
        "No factory towing approval.",
        "Average 3.8 service visits per year (battery-alert issues).",
        ’Owner survey: 41 %% "would buy again" (1 200 owners, Feb-25).’,
        "Not eligible for federal tax credit.",
        "OTA updates limited to infotainment."
Profiles were generated programmatically. Profiles included an agent’s name, age, city of residence, interests, planned use case for the car (e.g., commuting, road trips, etc.), and major pain points or concerns (e.g., fuel efficiency, winter driving, etc.).
E.2.2 Prompts
@@
    f"You start believing there’s a 50% chance the car is GOOD. "
@@
    f’JSON response: {{"message": "<message>"}}’
E.3 Validation of LLM Beliefs and Oracle
We now provide further experiments to validate the self-reported beliefs of the LLMs (as shown in figure 1(b)) and the performance of the LLM oracle. In these experiments, we utilize the LLMs as rational and utility maximizing, but not necessarily Bayesian, agents. The goal of these experiments is (1) to demonstrate that the model’s self-reported posterior genuinely corresponds to the belief that governs its choices in state-contingent decision problems, and (2) to verify that messages produced by the oracle do have the desired precision.
Figure 4: The rate at which the LLM chose a ‘buy’ decision as a function of (a) its self-reported posterior belief (b) its self-reported posterior and the price buying.
First, we provide the LLM with a prior and a new signal, obtain its self-reported posterior belief, and then ask it to make a binary “buy/not-buy” decision that yields unit reward if and only if the true state is good. A rational agent should choose “buy” exactly when its subjective belief exceeds 0.5. As shown in figure 4(a), the LLM’s decisions align sharply with this threshold: it overwhelmingly chooses “buy” when its reported posterior exceeds 0.5 and “not buy” otherwise. Next, we elicit the model’s willingness to pay for a payoff that is contingent on the true state. For a range of prices $c$, we ask whether the agent would pay $c$ in exchange for receiving unit utility if the state is good. A rational agent with belief $p$ should accept exactly when $c<p$. Figure 4(b) plots the fraction of “buy” decisions across posterior–price pairs, showing a clear separation: acceptance rates are high when $c<p$ and low when $c>p$, precisely as expected from a utility-maximizing decision maker whose belief is $p$. Both experiments demonstrate that the self reported posterior is close to the inherent belief of the model as used in decision making.
Similarly, we validate the performance of the oracle in crafting messages of a given precision. After the oracle generates a message of specified precision $q$, we ask another LLM to form an estimate $\hat{q}$ of the message’s precision. We do this for $q\in\{0.5,0.6,0.7,0.8,0.9,1.0\}$ on 100 programmatically generated agent profiles with 10 trials for each combination for a total of 6,000 trials. The mean absolute error |qq^||q-$\hat${q}| was just 0.025. This indicates that the realized precision of the oracles message’s does closely follow the requested precision.
E.4 Impact of Non-Bayesian LLM Belief Updates on Learning Trajectories